### v1

# AnyPcc: Compressing Any Point Cloud with a Single Universal Model

Abstract
Generalization remains a critical challenge for deep learning-based point cloud geometry compression. We argue this stems from two key limitations: the lack of robust context models and the inefficient handling of out-of-distribution (OOD) data. To address both, we introduce AnyPcc, a universal point cloud compression framework. AnyPcc first employs a Universal Context Model that leverages priors from both spatial and channel-wise grouping to capture robust contextual dependencies. Second, our novel Instance-Adaptive Fine-Tuning (IAFT) strategy tackles OOD data by synergizing explicit and implicit compression paradigms. It fine-tunes a small subset of network weights for each instance and incorporates them into the bitstream, where the marginal bit cost of the weights is dwarfed by the resulting savings in geometry compression. Extensive experiments on a benchmark of 15 diverse datasets confirm that AnyPcc sets a new state-of-the-art in point cloud compression. Our code and datasets will be released to encourage reproducible research.
Figure 1: (a) AnyPcc Architecture. A single, unified model compresses point clouds from any source, with our Instance-Adaptive Fine-Tuning (IAFT) module boosting performance on out-of-distribution (OOD) data. (b) AnyPcc Benchmark. Our comprehensive benchmark features 15 diverse datasets, including both standard and extreme cases. When compared against five state-of-the-art methods, AnyPcc consistently achieves high compression efficiency across all types of point clouds.
[…]
1 Introduction
Figure 2: The superiority of our UCM in capturing contextual information across diverse point cloud types (e.g., dense and sparse).
The proliferation of 3D content across applications like autonomous driving and virtual reality has established point clouds as a primary data representation. This ubiquity creates a critical need for efficient geometry compression algorithms to minimize storage and transmission costs. While recent learning-based methods surpass traditional standards like G-PCC [7], their real-world applicability is hampered by a significant generalization gap. We argue this gap originates from two fundamental limitations. First, existing context models are often tailored to specific point cloud densities and thus fail to maintain stable performance across the wide spectrum of densities found in real-world data, such as sparse LiDAR scans and dense reconstruction outputs. Second, these models exhibit a sharp decline in compression efficiency when encountering out-of-distribution (OOD) data (i.e., samples statistically dissimilar to their training corpus). We address these limitations sequentially, first by tackling context modeling.
Effective context modeling is pivotal for high-efficiency point cloud compression. For instance, SparsePCGC [49] establishes strong contextual priors through spatial auto-regressive grouping, but its performance degrades in sparse scenarios where local neighborhood information becomes unreliable. In contrast, RENO [64] re-frames the problem from spatial prediction to a channel-wise occupancy probability prediction task, thereby maintaining rich context even with sparse inputs. However, by relying exclusively on channel-wise relationships, it foregoes valuable spatial priors. This presents a fundamental dilemma: methods either leverage robust spatial context at the risk of failing in sparse conditions, or ensure sparsity-robustness by sacrificing spatial information, as shown in Figure 2. To resolve this trade-off, we propose a Universal Context Model (UCM) that synergistically integrates both spatial and intra-channel priors via a novel grouping strategy. This dual-pronged approach allows our model to capture a richer set of contextual dependencies, ensuring robust compression across the entire density spectrum.
While prior works [13, 41, 8] have explored Implicit Neural Representations (INRs) for generalization, their need to train a network from scratch for each instance leads to prohibitive encoding times, hindering their practicality. Conversely, while our pre-trained UCM provides a robust foundation, its performance can still degrade when faced with extreme OOD samples. To overcome both the slowness of INRs and the generalization limits of a fixed pre-trained model, we introduce AnyPcc. Our key insight is to synergize a general, powerful pre-trained model with rapid, instance-specific adaptation. Specifically, our Instance-Adaptive Fine-Tuning (IAFT) strategy achieves this by updating only a minuscule subset of the network’s parameters (the final linear layers), while keeping the vast majority frozen. This lightweight process, enabled by the strong priors from our UCM, allows AnyPcc to converge on a highly efficient, instance-specific model in mere seconds. Consequently, AnyPcc strikes an exceptional balance between compression performance, generalization, and practical efficiency, as shown in Figure 1. Our contributions are summarized as follows:
  • We introduce AnyPcc, a universal compression framework designed to resolve the critical trade-off between efficiency and generalization. To our knowledge, it is the first method to achieve both high compression rates and robust performance across diverse point cloud types using a single, unified model.
  • Our Universal Context Model (UCM) uses a novel checkerboard grouping strategy to synergistically integrate spatial and channel priors, establishing robust context modeling across all point cloud densities.
  • We pioneer an Instance-Adaptive Fine-Tuning (IAFT) strategy that resolves the trade-off between explicit and implicit compression. It rapidly fine-tunes a small subset of a pre-trained model’s weights for each instance, enabling high-fidelity compression in seconds.
  • Extensive experiments on a benchmark of 15 diverse datasets demonstrate that AnyPcc not only outperforms the latest G-PCC v23 standard but also sets a new state-of-the-art among learning-based methods, establishing a new paradigm for universal point cloud compression.
2 Related Work
2.1 Point Cloud Geometry Compression
2.1.1 Category-Specific Methods.
In the domain of object point cloud compression [59, 63, 39, 60, 61, 35], several methods leverage sparse tensor representations. For instance, PCGCv2 [48], SparsePCGC [49], UniPCGC [52] and methods [28, 68, 66, 5, 38, 71, 15] employ sparse convolutions to exploit local correlations for effective compression. Concurrently, approaches like VoxelDNN [36] and SparseVoxelDNN [34] achieve higher compression efficiency through autoregressive codecs for feature extraction. For LiDAR point clouds compression [65, 44, 10, 40, 9, 3, 31, 58, 25], seminal works such as MuSCLE [6], Octsqueeze [22], OctAttention [16], EHEM [43] and TopNet [56] utilize octree-based [42] representations, employing attention mechanisms to predict the occupancy codes of parent nodes efficiently. Other methods, including SparsePCGC and RENO [64], also adapt the sparse tensor paradigm to LiDAR data. More recently, with the rise of 3D Gaussian Splatting [26, 30] (3DGS), GausPcgc [53] is introduced, which enables efficient compression of 3DGS coordinates. However, a common thread unites these diverse methods: while effective, their architectures are highly tailored to the unique characteristics of their target point cloud domain, limiting their universal applicability.
2.1.2 The Generalization Problem.
While recent efforts have aimed for versatility, exemplified by frameworks like Unicorn [50] and standardization initiatives (MPEG [1, 2], AVS, JPEG [18] AI-PCC), their underlying methodologies remain constrained by category-specific architectures and dedicated training datasets. A prime example is Unicorn-U, which, despite its goal of unification, reveals two critical flaws. First, it employs a non-unified, hybrid architecture where distinct feature extractors (e.g., attention and convolution) are manually prescribed for different sample types, compromising its practical utility. Second, and more fundamentally, its performance is confined to the data distributions seen during training, exhibiting a sharp collapse when encountering out-of-distribution (OOD) data types. This dependency on curated training data is untenable for the vast and varied landscape of real-world point clouds. Dedicated training data is often unavailable for many critical types, such as medical scans, 3D Gaussian Splats, or point clouds generated by networks like Dust3R [55] and VGGT [51]. Consequently, the compression efficiency of existing methods degrades drastically on such data. This performance collapse reveals a critical, yet largely unaddressed, challenge in the current research landscape: the true generalization capability of point cloud compression networks.
2.2 Implicit Compression
Figure 3: llustration of the proposed AnyPcc framework. AnyPcc formulates the compression task as a sequential prediction of multi-scale occupancy codes. It develops a Universal Context Model (UCM) that incorporates strong inductive biases from spatio-channel partitioning, enabling robust generalization across diverse data sources.
Implicit neural representations (INRs) have emerged as a potent solution to the generalization challenge. This paradigm involves overfitting a coordinate-based network, typically a Multi-Layer Perceptron (MLP), to a single data instance [13, 69, 14, 8, 41, 62, 47, 33, 70]. The optimized network weights themselves form the compressed representation. While this per-instance optimization grants excellent generalization, it is hampered by critical drawbacks: the encoding process is computationally prohibitive as it equates to full network training, and the fitting procedure is lossy.
Some works have sought to mitigate these issues. For instance, LINR-PCGC [23] introduced implicit representations into lossless compression by fine-tuning a network with shared weights across multiple frames. However, this approach still requires tuning the entire network and is limited to multi-frame scenarios. In the domains of image and video compression [32, 46, 37], parameter-efficient fine-tuning (PEFT)  [54, 21] techniques, such as adapters, have shown success in achieving strong generalization by adapting pre-trained models. Yet, such a hybrid paradigm, combining the strengths of explicit priors and implicit adaptation, remains unexplored in point cloud compression. Furthermore, the highly irregular distribution of point clouds necessitates a tailored investigation into how these explicit-implicit models can be effectively designed.
3 Method
3.1 Overview
Our proposed framework, AnyPcc, achieves universal point cloud compression through three core components. First, we introduce a Universal Context Model (UCM), pre-trained on diverse data to adaptively handle point clouds of varying densities. Second, we employ an Instance-Adaptive Fine-tuning strategy to efficiently compress out-of-distribution (OOD) samples. Finally, we demonstrate how our framework can be seamlessly extended from lossless to lossy compression via a probability thresholding mechanism, creating a single, unified solution. These components are detailed in the following subsections.
3.2 Universal Context Model
Design Insight.
Our Universal Context Model (UCM) is designed to overcome a fundamental limitation in prior art: the trade-off between performance on dense versus sparse point clouds. Spatial context models [49] excel on dense data but falter in sparse regions, whereas channel-wise models [64] are sparsity-robust but underutilize inter-voxel dependencies in dense areas. To resolve this, the UCM introduces a synergistic fusion of spatial and channel-wise context partitioning, creating a model that is robust across the entire density spectrum, as shown in Figure 2 and 3.
This spatio-channel design is grounded in two theoretical principles, which we prove in Appx.. We first establish that modeling on 8-bit occupancy code channels is information-theoretically equivalent to modeling on the corresponding $2\times 2\times 2$ voxel block (Theorem 1). We then prove that this approach provides a significant receptive field advantage: a convolution with a kernel of size $k$ on the occupancy codes is equivalent to using a kernel of size $2k$ on the voxel space, which is crucial for capturing context in sparse data (Theorem 2).
Leveraging these principles, the UCM’s coarse-to-fine framework is two-fold: it operates across a multi-scale hierarchy for prediction and also follows a coarse-to-fine predictive path (intra-scale) within each level. Critically, this process must be distinguished from image context models [19, 20, 24]. Although the context grouping appears similar, it fundamentally operates on geometric occupancy codes, thereby creating a coarse-to-fine geometric partitioning rather than modeling a latent space. This dual design ensures robust, density-adaptive prediction across all conditions, as detailed in Figure 4. (More in Appx.)
Notation.
We define the symbols used as follows. V2O (Voxel-to-Occupancy-Code) and O2V (Occupancy-Code-to-Voxel) denote the conversion processes illustrated in Figure 2. Oc-Code stands for occupancy code, and OV for occupied voxels. AE / AD (Arithmetic Encoder / Decoder) are responsible for compressing the occupancy codes into a bitstream based on their predicted probabilities.
Hierarchical Context Propagation.
Our model adopts a coarse-to-fine hierarchical structure. At each level $l$, the geometric information is captured by a set of occupied voxel coordinates $\mathcal{C}^{(l)}$ and their associated 8-bit occupancy codes $\{o_{i}^{(l)}\}$. The context propagation mechanism, involves two main stages. First, a coarse-context encoding network, $\Psi_{\text{prior}}$, processes the occupancy codes and coordinates from level $l$ to generate a powerful latent representation $\mathbf{Z}^{(l)}$. This network implicitly handles the conversion from discrete codes to continuous features and aggregates local contextual information:
$\mathbf{Z}^{(l)}=\Psi_{\text{prior}}(\{o_{i}^{(l)}\},\mathcal{C}^{(l)}).$ (1)
Second, this latent representation is propagated to the finer resolution level $l+1$. The features are upsampled and then refined by a target-processing network $\Psi_{\text{target}}$ to produce the final predictive context, $\{\mathbf{c}_{j}^{(l+1)}\}$, for each voxel at the finer scale:
$\{\mathbf{c}_{j}^{(l+1)}\}=\Psi_{\text{target}}(\text{Upsample}(\mathbf{Z}^{(l)})),$ (2)
where $\Psi_{\text{target}}$ and $\Psi_{\text{prior}}$ are shown in Figure 4. This resulting context field, $\{\mathbf{c}_{j}^{(l+1)}\}$, provides the foundation for our detailed spatio-channel prediction task.
Spatio-Channel Context Factorization.
We decompose the joint probability distribution of the occupancy codes at a given level by factorizing the context along both spatial and channel-wise dimensions. This dual factorization dramatically reduces the complexity of the predictive task and improves coding efficiency.
Spatial Context Partitioning.
We first partition the target voxels at level $l+1$ into two disjoint, non-adjacent sets using a 3D checkerboard pattern, as illustrated in Figure 3. A voxel with coordinates $(x,y,z)$ is assigned to the first group, $\mathcal{G}_{1}$ (yellow cells), if its coordinate sum is even, and to the second group, $\mathcal{G}_{2}$ (green cells), otherwise:
$\mathcal{G}_{1}$ $=\{i\in\mathcal{V}^{(l+1)}\mid(x_{i}+y_{i}+z_{i})\pmod{2}=0\},$ (3)
$\mathcal{G}_{2}$ $=\{i\in\mathcal{V}^{(l+1)}\mid(x_{i}+y_{i}+z_{i})\pmod{2}=1\}.$ (4)
This partitioning transforms the prediction task into a two-step auto-regressive process. The joint probability of all occupancy codes $\{o_{i}\}$ at the current level is factorized as:
$P(\{o_{i}\})=\prod_{i\in\mathcal{G}_{1}}P(o_{i}|\mathbf{c}_{i})\cdot\prod_{j\in\mathcal{G}_{2}}P(o_{j}|\mathbf{c}_{j},\{o_{k}\}_{k\in\mathcal{N}(j)\cap\mathcal{G}_{1}}),$ (5)
where $\mathbf{c}_{i}$ and $\mathbf{c}_{j}$ are the initial contexts from the coarser layer, and $\mathcal{N}(j)$ is the $n^{3}$-connectivity neighborhood of voxel $j$. This spatial grouping establishes a powerful local dependency model, allowing the network to leverage information from immediately adjacent, already-decoded voxels for more accurate probability estimation.
Figure 4: Our Universal Context Model (UCM) uses a recursive, parameter-shared architecture. At each scale $l$, the model takes the voxels of that scale as input to predict their corresponding occupancy codes, repeating this process with a consistent structure across all scales, as illustrated in the transition from scale $l$ to $l+1$.
Channel Context Partitioning.
Within each spatial group, we further decompose the 8-bit occupancy code $o\in\{0,\dots,255\}$ into two 4-bit sub-symbols, a least significant part $o_{0}$ and a most significant part $o_{1}$.
$o_{0}=o\pmod{16},\quad o_{1}=\lfloor o/16\rfloor.$ (6)
This channel-wise partitioning enables a two-stage prediction cascade, factorizing the probability of the occupancy code for any voxel $i$ as:
$P(o_{i}|\mathbf{c}_{i})=P(o_{i,0}|\mathbf{c}_{i})\cdot P(o_{i,1}|o_{i,0},\mathbf{c}_{i}),$ (7)
where $\mathbf{c}_{i}$ is the feature vector representing the context for a voxel $i\in\mathcal{G}_{1}$. The probabilities of the sub-symbols are modeled sequentially. For the first group $\mathcal{G}_{1}$, this is formulated as:
$P(o_{i,0}|\mathbf{c}_{i})$ $\propto\exp\left(\text{MLP}_{\text{g1,0}}\left(\Phi_{\text{g1,0}}(\mathbf{c}_{i})\right)\right),$ (8)
$P(o_{i,1}|o_{i,0},\mathbf{c}_{i})$ $\propto\exp\left(\text{MLP}_{\text{g1,1}}\left(\Phi_{\text{g1,1}}\left(\mathbf{c}_{i}+\text{Emb}(o_{i,0})\right)\right)\right).$ (9)
This cascading mechanism, which uses dedicated prediction heads ($\text{MLP}_{\text{g1,0}},\text{MLP}_{\text{g1,1}}$) for the first group, adaptively refines the context and significantly improves coding efficiency. A similar process is applied to the $\mathcal{G}_{2}$.
Synergistic Feature Aggregation.
The synergy between spatial and channel partitioning is most evident in the encoding of the second spatial group, $\mathcal{G}_{2}$. The context vector for a voxel $j\in\mathcal{G}_{2}$ is explicitly enhanced by aggregating features from its decoded neighbors in $\mathcal{G}_{1}$. This process is highlighted by the red dotted circle in Figure 3.
Let $\mathbf{F}$ denote the latent feature tensor produced by $\Psi_{\text{target}}$, where each voxel location $k$ is associated with a latent feature vector $\mathbf{f}_{k}$. To condition the encoding of $\mathcal{G}_{2}$ on the outcomes of $\mathcal{G}_{1}$, we first construct a sparse, augmented feature tensor $\hat{\mathbf{F}}$. This tensor is populated exclusively at the voxel locations corresponding to the already-decoded group $\mathcal{G}_{1}$. Specifically, for each voxel $k\in\mathcal{G}_{1}$, we augment its feature vector $\mathbf{f}_{k}$ with an embedding of its occupancy symbol $o_{k}$:
$\hat{\mathbf{f}}_{k}=\mathbf{f}_{k}+\text{Emb}_{\text{prior}}(o_{k}),\quad\forall k\in\mathcal{G}_{1}.$ (10)
Next, we gather this local context for the subsequent group using a sparse convolution. The aggregated feature $\mathbf{f}_{\text{agg},j}$ for each voxel $j\in\mathcal{G}_{2}$ is computed by summing the features of its neighbors in $\mathcal{G}_{1}$, weighted by the kernel $\mathbf{W}_{\text{sconv}}$:
$\mathbf{f}_{\text{agg},j}=\left[\text{Sconv}(\hat{\mathbf{F}})\right]_{j}=\sum_{k\in\mathcal{N}(j)\cap\mathcal{G}_{1}}\mathbf{W}_{\text{sconv}}(k-j)\cdot\hat{\mathbf{f}}_{k}.$ (11)
This operation efficiently collects decoded geometric information from the spatial neighborhood $\mathcal{N}(j)$. The final, enhanced context feature $\mathbf{c}^{\prime\prime}_{j}$ for voxel $j$ is then produced by passing the original voxel feature $\mathbf{f}_{j}$ along with its aggregated neighbor feature $\mathbf{f}_{\text{agg},j}$ through a fusion network $\Phi_{\text{fuse}}$:
$\mathbf{c}^{\prime\prime}_{j}=\Phi_{\text{fuse}}([\mathbf{f}_{j},\mathbf{f}_{\text{agg},j}]),$ (12)
where $\cdot$$\cdot$] denotes concatenation. This process allows the model to directly leverage the decoded geometric structure from $\mathcal{G}_{1}$ to resolve ambiguities and improve predictions for the subsequent group $\mathcal{G}_{2}$. Finally, an arithmetic encoder utilizes the predicted probabilities to compress the ground-truth occupancy codes into a bitstream. On the decoding side, an arithmetic decoder employs the same probabilities to reconstruct the occupancy codes from the bitstream.
3.3 Instance-Adaptive Fine-Tuning
While the pre-trained UCM demonstrates high efficiency across a wide range of point clouds, its performance can degrade when encountering test instances with statistical distributions that significantly deviate from the training datasets—a common challenge known as out-of-distribution (OOD) generalization. To bridge this generalization gap, we introduce a novel instance-adaptive fine-tuning strategy. This approach leverages the principles of implicit neural representations (INRs) to create a powerful hybrid coding paradigm, ensuring robust performance even on previously unseen point cloud types.
The core idea is to make the network weights themselves a compact representation of a specific point cloud’s geometry. To achieve this, we partition the UCM’s parameters $\Theta$ into a large, frozen backbone $\Theta_{\text{frozen}}$ and a small set of tunable parameters $\Theta_{\text{tune}}$, such that $\Theta=\Theta_{\text{frozen}}\cup\Theta_{\text{tune}}$. The frozen set $\Theta_{\text{frozen}}$ contains the feature extraction and spatial convolution modules, which capture generalizable priors. The tunable set $\Theta_{\text{tune}}$ consists exclusively of the lightweight parameters within the final prediction heads ($\text{MLP}_{\text{g1,0}},\text{MLP}_{\text{g1,1}},\text{MLP}_{\text{g2,0}}$ and $\text{MLP}_{\text{g2,1}}$). In practice, this fine-tuning is highly efficient. We first perform a single forward pass through the frozen backbone ($\Theta_{\text{frozen}}$) to compute and cache the intermediate features. The subsequent training iterations then operate solely on these cached features to update the lightweight prediction heads ($\Theta_{\text{tune}}$), dramatically accelerating the process.
For each new point cloud $P$, we perform a rapid on-the-fly optimization to find a specialized set of weights $\Theta_{\text{tune}}^{*}$ that minimizes an instance-specific loss function. This optimization process is formulated as:
$\Theta_{\text{tune}}^{*}=\arg\min_{\Theta_{\text{tune}}}\mathcal{L}_{\text{tune}}(P;\Theta_{\text{tune}}).$ (13)
The loss function $\mathcal{L}_{\text{tune}}$ is composed of the instance’s bitrate, modeled by the negative log-likelihood of its occupancy symbols $\{o_{i}\}$, and an L1 regularization term to promote parameter sparsity:
$\mathcal{L}_{\text{tune}}$ $=\mathcal{L}_{\text{instance}}+\lambda_{\text{L1}}\|\Theta_{\text{tune}}\|_{1},$ (14)
$\text{where}\quad\mathcal{L}_{\text{instance}}$ $=-\sum_{i=1}^{|P|}\log p(o_{i}|\text{context}_{i};\Theta).$ (15)
Because $\Theta_{\text{tune}}$ is small, this process converges in just a few hundred gradient steps, turning the optimized weights $\Theta_{\text{tune}}^{*}$ into a highly accurate, instance-specific generative model.
The final compressed bitstream $\mathcal{B}(P)$ for the point cloud is then constructed from two components:
  1. 1.
    The Implicit Model Component ($\mathcal{B}_{\text{weights}}$). The fine-tuned weights $\Theta_{\text{tune}}^{*}$ are quantized to $\hat{\Theta}_{\text{tune}}^{*}$ and then losslessly compressed using ‘deepCABAC‘ [57], a specialized arithmetic coder for neural network parameters.
  2. 2.
    The Geometry Component ($\mathcal{B}_{\text{geom}}$). The occupancy symbols $\{o_{i}\}$ are losslessly compressed using an arithmetic coder guided by the instance-adapted probability distributions $p(\cdot|\hat{\Theta}_{\text{tune}}^{*},\Theta_{\text{frozen}})$.
The total bitrate is $R(P)=|\mathcal{B}_{\text{weights}}|+|\mathcal{B}_{\text{geom}}|$, where $|\cdot|$ denotes the length of the bitstream.
At the decoder, the process is mirrored: it first decodes $\mathcal{B}_{\text{weights}}$ to reconstruct the specialized model parameters $\hat{\Theta}_{\text{tune}}^{*}$, and then uses this instance-specific model to entropy-decode $\mathcal{B}_{\text{geom}}$. By overfitting a small part of the network, we create a highly accurate probability model. The cost of transmitting the model, $|\mathcal{B}_{\text{weights}}|$, is significantly outweighed by the bitrate reduction for encoding the geometry, $|\mathcal{B}_{\text{geom}}|$, leading to substantial gains in compression efficiency and generalization.
3.4 Unified Lossless and Lossy Compression
Although presented primarily for lossless compression, our UCM framework seamlessly extends to lossy scenarios. While a naive approach of simply omitting entropy coding at certain scales is possible, it can cause severe geometric degradation in dense point clouds. We therefore adopt a more sophisticated strategy: the encoder transmits only the ground-truth point count ($k$) for a given scale. The decoder then reconstructs the geometry by identifying the $k$ most probable occupied locations and their corresponding occupancy codes based on the model’s predictions. For a comprehensive description of this method, see the Appx..
4 Experimental Results
4.1 Experiment Setup
Dataset.
To cultivate a powerful and generalizable Universal Context Model (UCM), we curate a comprehensive training corpus by merging multiple datasets. This corpus includes a wide array of point clouds recommended by the MPEG AI-PCC and AVS AI-PCC working groups, such as KiTTI [17], Ford [4], 8iVFB [12], MVUB [29], and ScanNet [11]. To further diversify the training data and enhance model robustness, we also incorporate the GausPcc-1K [53, 27] and Thuman [67] datasets. To truly challenge our proposed framework and assess its real-world viability, our evaluation protocol goes far beyond standard benchmarks. In addition to all mainstream datasets, we deliberately introduce point clouds from modern reconstruction techniques like VGGT methods and 3D Gaussian Splatting. Furthermore, we synthesize three challenging datasets to simulate common data imperfections: NS (added noise), RS (point dropout), and CS (non-rigid shape deformations). This comprehensive and demanding benchmark provides a more realistic measure of a model’s stability and generalization capabilities (see Appx.).
Table 1: Performance comparison on the AnyPcc Benchmark. The table presents the compression performance of AnyPcc against six methods across 15 diverse datasets, with the best and second-best results highlighted in red and yellow cells.
Dataset Cond OOD RENO SparsePCGC Unicorn* OctAttention TopNet GPCC Ours Ours-U
8iVFB E 0.70 0.57 0.57 0.68 0.59 0.76 0.54 0.57
MVUB 1.00 0.69 0.69 0.76 0.69 0.94 0.67 0.75
Owlii 0.59 0.48 0.48 0.62 0.56 0.59 0.47 0.47
Thuman 1.64 1.70 1.70 2.31 2.20 2.00 1.58 1.64
ScanNet 2.15 1.86 1.86 2.13 2.03 2.03 1.83 1.88
KITTI 7.06 6.80 6.50 7.21 6.85 8.19 6.18 6.45
Ford 9.38 9.77 8.44 9.10 8.54 10.32 8.40 8.57
Dense M 5.81 6.37 5.48 6.55 6.38 5.32 5.27 5.55
Sparse 9.64 9.98 9.42 10.40 10.02 9.35 9.11 9.26
GS 13.89 15.82 / 11.31 10.95 14.46 11.65 11.74
VGGT 8.24 7.84 / 8.22 7.83 7.33 7.30 7.06
S3DIS 13.06 11.88 / 11.52 10.84 10.66 10.93 10.79
RS H 4.02 3.88 / 4.05 3.92 3.72 3.68 3.50
NS 4.96 6.54 / 4.89 4.84 4.85 4.69 4.67
CS 3.94 4.94 / 3.40 3.21 3.23 3.18 3.08
CR Gain over GPCC $\downarrow$ 2.96% 2.07% / 1.32% -4.04% 0.00% -11.93% -10.75%
Enc/Dec Time (s) $\downarrow$ 0.22/0.23 2.6/2.2 / 7.7/1324 8.7/1740 3.8/2.7 2.84/0.46
Total Parameters (M) $\downarrow$ 9.03 26.43 / 29.61 23.59 / 68.39 9.77
  • Cond represents the test difficulty of the testsets, and we divide the test set into easy (E), medium (M), and hard (H).
    The results for Unicorn are cited directly from the original publication as its implementation is not open-source.
Implementation.
Our framework is built with PyTorch and TorchSparse [45], and all experiments are conducted on a single NVIDIA RTX 3090 GPU. We propose and evaluate two distinct versions of our model:
  • Ours: This version aligns with standard practice by training a dedicated model for each dataset category.
  • Ours-U: This is a single, unified model trained on a large-scale mixed dataset. The same set of weights is applied to all test sets, greatly improving its practical utility.
For both versions, we apply our UCM on E samples, and augment it with IAFT (200 iters) for M and H samples.
Baseline and Metrics.
We benchmark our method against several state-of-the-art (SOTA) open-source solutions, including RENO [64], SparsePCGC [49], OctAttention [16], and TopNet [56], as well as the latest traditional codec, GPCC v23. Since the code for Unicorn [50] is unavailable, we report its performance from the original paper under aligned experimental settings. All models are evaluated under identical training and testing conditions for a fair comparison. For lossless compression, we report the efficiency in bits per point (bpp) and the Compression Ratio Gain (CR-Gain). The CR-Gain is calculated relative to an anchor codec as $(\text{bpp}_{\text{method}}-\text{bpp}_{\text{anchor}})/\text{bpp}_{\text{anchor}}\times 100\%$, where a more negative value indicates greater bitrate savings. For lossy compression, we assess the rate-distortion performance using bpp for the rate and Peak Signal-to-Noise Ratio (PSNR) for the distortion.
4.2 Lossless Compression
Results.
As benchmarked in Table 1 across 15 datasets (including 5 for Out-of-Distribution generalization), our methods demonstrate clear superiority. Our models, Ours and Ours-U, achieve SOTA on 13 datasets and deliver significant BD-Rate savings of 11.93% and 10.75% over the GPCC v23 anchor. In stark contrast, most baselines fail to match this anchor; RENO, SparsePCGC, and Octattention show performance degradations of 2.96%, 2.07%, and 1.32%. Among the baselines, only TopNet provides a positive gain of 4.04%. A key trade-off emerges: our specialized model excels on in-distribution data, but our universal model (Ours-U) shows superior generalization across all OOD datasets, affirming its practical value. For OOD evaluation, models trained on KITTI are used. Moreover, we provide both the results for RENO-U and SparsePCGC-U (Unicorn-U) with unified data training and the evaluations on dense point clouds in Appx..
Figure 5: Performance comparison using rate-distortion curves. Comparisons on the human-body and ScanNet datasets exclude RENO, OctAttention, and TopNet, as they are specifically designed for lossy LiDAR compression and thus incompatible.
Model Parameters.
Our primary method (Ours), like all baselines, requires training seven distinct models for the benchmark, as shown in Table 1. In contrast, our universal model (Ours-U) utilizes a single set of weights for all datasets. This not only highlights its generalization but also drastically reduces storage and deployment overhead, marking a significant step towards a practical solution.
Codec Time.
Our decoding time is comparable to RENO, the fastest baseline, as shown in Table 1. While our default encoding takes 2.84 seconds, it is highly flexible. By adjusting fine-tuning iterations, the encoding time can be controlled within a 0.44s to 2.84s range, offering a trade-off between speed and compression efficiency. A detailed analysis is in the Appx..
4.3 Lossy Compression
The lossy compression results are shown in Figure 5. LiDAR-based methods are not shown for the dense human-body and ScanNet datasets due to their suboptimal performance. The results confirm that AnyPcc performs robustly across all datasets. Crucially, the strong performance of our unified model, Ours-U, demonstrates that AnyPcc can serve as a single-model solution for both efficient lossless and lossy point cloud compression. For an analysis of model parameters and codec times in the lossy compression, please refer to the Appx..
4.4 Ablation Studies
Table 2: Effectiveness analysis of key modules in UCM.
SC SG CG CR-Gain Params (M)
Baseline 0.00% 5.15
Abla1 -6.56% 5.68
Abla2 0.13% 5.15
Abla3 -7.74% 9.78
Abla4 -5.33% 7.19
Abla5 -6.50% 5.67
Ours -9.88% 9.77
Key Modules.
Our ablation study on the key modules of Spatial Convolution (SC), Spatial Grouping (SG), and Channel Grouping (CG) is presented in Table 2. The results confirm that utilizing all three components simultaneously achieves the best performance. Notably, while SC and SG individually improve performance, using only CG, as is the case in RENO, actually degrades it. This indicates that channel priors from CG are only effective when coupled with the strong spatial priors provided by SC and SG.
Channel Count.
Table 3 shows the effect of varying the channel count (32/64/128) on performance and complexity. While higher channel counts lead to better performance, the model with 128 channels proves to be too computationally expensive. We therefore choose C=64 for our model, as it offers the best trade-off between performance and speed.
IAFT Module.
We also conduct an ablation study on the Instance-Adaptive Fine-Tuning (IAFT) module, with the results presented in Table 3. The table shows that fine-tuning for 800 iterations yields a performance gain of approximately 5%, at the cost of an additional 10 seconds in encoding time. This allows for dynamic control over the complexity-performance trade-off. As shown in Table 4, IAFT reduces the entropy-coded bpp by 1.883 on the GS dataset, at the cost of 0.319 bpp for the model. This confirms that for samples with irregular distributions, the bitrate savings from entropy coding substantially outweigh the network overhead.
Table 3: Ablation study of channel count and IAFT module.
UCM UCM+IAFT
Bpp 5.42/5.32/5.25 5.14/5.04/4.99
Enc Time (s) 0.31/0.44/1.13 11.82/12.11/13.08
Dec Time (s) 0.32/0.46/1.14 0.32/0.46/1.14
Model Size (M) 2.45/9.77/39.01
Table 4: Ablation study of IAFT on bpp composition (GS dataset).
Bpp Entropy Coding Net Size
UCM 13.307 0
UCM+IAFT 11.424 0.319
More Ablation Studies.
Ablation studies for the IAFT module, the training data ratio, and the context model design are provided in the Appx..
5 Conclusion
This paper introduces AnyPcc, a framework for geometry point cloud compression. Powered by a robust generic context model and an instance-adaptive fine-tuning strategy, our approach uniquely combines the advantages of both explicit and implicit compression methods. This allows for effective compression of both in-distribution (ID) and out-of-distribution (OOD) samples. Experimental results on a comprehensive benchmark of 15 datasets show that AnyPcc achieves state-of-the-art compression performance.